On November 1, 20-n, you notice the following bid-ask quotes in the option market (supposedly perfect otherwise!), the underlying asset being stock ABC:

Call 130 January: 11-12

Call 145 January: 7-8

Put 130 January: 8.5-9.5

Put 145 January: 16-17

ABC quotes 133 and the 3-month interest rate is 4% (proportional, annualized rate).

a) What arbitrage should you undertake (transaction costs are negligible, but you are not market-maker, so that you are subject to the bid-ask spread)?

b) Is it really riskless?

c) What is its IRR if you do not cancel, either by lending or by borrowing, the initial cash-inflow or outflow?
Answer #1

Every option has two prices at any time of the trading day. The first price is called the bid or sell price, and its the p

Recall the Call Put Parity Equation:

C - P = S - PV (K)

K = strike Price = 130

LHS = C - P = Buy a call + Short (Sell) a Put = 12 - 8.5 = 3.5

RHS = S - PV (K) = 133 - 130 x (1 + 4% x 3 / 12)-1 = 4.29

RHS = S - PV(K) > LHS = C - P

Hence, S - PV(K) + P - C > 0

Part (a)

Hence, the arbitrage strategy should be:

  • Borrow Present value of the strike price
  • Short sell the call
  • Buy the stock
  • Buy the put option

Part (b)

Yes, it is really riskless.

Part (c)

IRR = Riskless return = 4%